OpenStudy (anonymous):

What is the slope of a line that passes through the point (−5, 3) and is parallel to a line that passes through (2, 13) and (−4, −11)?

OpenStudy (anonymous):

Can you find the slope of the line that passes through (2, 13) and (−4, −11)?

OpenStudy (anonymous):

thats the thing i dont know how to. i dont have my notes with me

OpenStudy (anonymous):

This is essential knowledge that you need to memorize as long as you want to continue your math study: slope m = Δy/ Δx

OpenStudy (anonymous):

Can you calculate now?

OpenStudy (anonymous):

so m= 13y/ -4x ?

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

I mean: Δy = 13 - ( -11 ) = .. Δx = ...

OpenStudy (anonymous):

x= 24

OpenStudy (anonymous):

NO, Δy = 24 ( not, x = 24) Now similar Δx = ...

OpenStudy (anonymous):

im sorry with math im slow. so its y=24. but what do you mean now similar ?

OpenStudy (anonymous):

Similar means use Δ to find the difference ( subtraction): Δy = y1 - y2 = 24 Δx = x1 - x1 = ...?

OpenStudy (anonymous):

x= 6

OpenStudy (anonymous):

Yes, m = Δy/ Δx = ....?

OpenStudy (anonymous):

/ symbol gets me confused, do i mult. or divide.

OpenStudy (anonymous):

so i got =6

OpenStudy (anonymous):

again, the distance Δx = 6

OpenStudy (anonymous):

Now slope means the fraction between Δy / Δx = ....?

OpenStudy (anonymous):

but i divided the x &&y so how can i fraction them out?

OpenStudy (anonymous):

24?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

x= 6*

OpenStudy (anonymous):

Then m = Δy / Δx = ... ?

OpenStudy (anonymous):

lol. soo the slope is 4? because i divided the y&&x

OpenStudy (anonymous):

y= 24 x= 6

OpenStudy (anonymous):

AGAIN, m = Δy / Δx = 24/ 6 = 4 You will be confused forever if you don't learn the right notation from now on!